A. Solve the following problems:

1. On the assumption that variables are continuous, compute for the median of the following data:

a. 1, 4, 5, 9, 11

b. 4, 8, 16, 27, 52

c. 7, 26, 31, 36, 46, 65

2. Calculate the range for the following sets of measurements:

a. 3, 6, 7, 10, 2

b. 11, 27, 32, 45, 54

3. Calculate the variance and standard deviation of the following: 13, 18, 21, 19, 16, 15.



1
Expert's answer
2022-01-14T12:15:00-0500

1.

a. Median "\tilde{x}=5," third value

b. Median "\tilde{x}=16," third value

c. Median "\tilde{x}=33.5, \dfrac{31+36}{2}=33.5"


2.

a. "2, 3, 6, 7, 10"


"Range=10-2=8"

b. "11, 27, 32, 45, 54"


"54-11=43"

3.


"mean=\bar{x}=\dfrac{1}{6}(13+18+21+19+16+15)"

"=17"

"Var(X)=\sigma^2 =\dfrac{1}{6}((13-17)^2+(18-17)^2"

"+(21-17)^2+(19-17)^2+(16-17)^2"

"+(15-17)^2)=7"

"\sigma=\sqrt{\sigma^2}=\sqrt{7}\approx2.645751"